Find the general solution of the ordinary differential equation (D2+2D+1)y=e^-xlnx using method of variation of parameters.
1
Expert's answer
2021-08-02T10:52:15-0400

Let us find the general solution of the ordinary differential equation "(D^2+2D+1)y=e^{-x}\ln x" using method of variation of parameters. The characteristic equation "k^2+2k+1=0" of the homogeneous differential equation is equivalent to "(k+1)^2=0," and hence has the roots "k_1=k_2=-1." The general solution of the non-homogeneous equation is of the form "y=C_1(x)e^{-x}+C_2(x)xe^{-x}."


Therefore, we should solve the following system


"\begin{cases}C_1'e^{-x}+C_2'xe^{-x}=0\\-C_1'e^{-x}+C_2'e^{-x}-C_2'xe^{-x}=e^{-x}\ln x
\end{cases}"


which is equivalent to


"\begin{cases}C_1'=-C_2'x\\-C_1'+C_2'-C_2'x=\ln x
\end{cases}"


"\begin{cases}C_1'=-C_2'x\\C_2'x+C_2'-C_2'x=\ln x
\end{cases}"


"\begin{cases}C_1'=-C_2'x\\C_2'=\ln x
\end{cases}"


"\begin{cases}C_1'=-x\ln x\\C_2'=\ln x
\end{cases}"


It follows that


"C_1(x)=-\int x\ln xdx="


|"u=\ln x, dv =xdx, du=\frac{dx}{x},v=\frac{x^2}{2}" |


"=-\frac{1}{2}x^2\ln x+\frac{1}{2}\int xdx"


"=-\frac{1}{2}x^2\ln x+\frac{x^2}{4}+C_1"


"C_2(x)=\int\ln xdx="



|"u=\ln x, dv =dx,du=\frac{dx}{x}, v=x" |


"=x\ln x-\int dx"


"=x\ln x-x+C_2."


We conclude that the general solution is


"y=(-\frac{1}{2}x^2\ln x+\frac{x^2}{4}+C_1)e^{-x}+(x\ln x-x+C_2)xe^{-x}" or


"y=(\frac{1}{2}x^2\ln x-\frac{3}{4}x^2+C_1+C_2x)e^{-x}."